OpenStudy (anonymous):

solve for x 4 - 5/(3-x) = 2/(x^2-9)

OpenStudy (saifoo.khan):

\[4 - \frac{5}{3-x} = \frac{2}{x^2-9}\]this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please help

OpenStudy (saifoo.khan):

Take LCD.

OpenStudy (anonymous):

ok